Database last updated on Mon, 02 Feb 2026 01:10:59 +0000 by localhost
Critical warnings from b-files' comments for days between 2024114 and 2024114 (1 sec, 4 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| hp: | Spectrometer alignment trouble. | 2 | 1 | | | #021 | 11:20:54 | | 1 | -1 |
| hg: | Lamp not on... hg test terminated F1= 0 F4= 206 | 4 | 1 | | | #021 | 11:24:00 | | 1 | -1 |
| hg: | Lamp not detected f1= 0 f4= 206 | 1 | 1 | | | #021 | 11:24:01 | | 1 | -1 |
| ze: | Zenith ZEROING FAILURE | 2 | 2 | | | #021 | 17:27:13 | | 1 | -1 |
Instrument constant changes for the week before 2024114 (1 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 3 rows)
Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)
| Serial | YYYYJJJ | HG large | max correction |
| 021 | 2024114 | 8 | 6.6 |
Non-critical warnings from b-files' comments for days between 2024114 and 2024114 (0 sec, 1 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| pd: | Going to menu | 1 | 99 | | | #021 | 17:27:31 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | dt | hg | hp | rs | sl |
| 114 | TORONTO | #021 | Attempting recovery FR,Must do HG now | 5 | 17 | 10 | 5 | 6 |
(0 sec)
Relative and absolute humidity for day 2024114 (1 sec, 1 rows)
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
| 021 | 3.8 | 1.9 |